Melissa is 24 years younger than Joyce. In 2 years Joyce will be 3 times as old as Melissa will be then. How old are each of them now?
Answer by elima(1433)

Melissa is 24 years younger than Joyce. In 2 years Joyce will be 3 times as old as Melissa will be then. How old are each of them now?
M=J-24
J=3M+2
J=3(J-24)+2
J=3J-72+2
J=3J-70
J-3J=-70
-2j=-70
J=35
M=35-24
M=11
:)